### v1

# On Trojans in Refined Language Models

Abstract
A Trojan in a language model can be inserted when the model is refined for a particular application such as determining the sentiment of product reviews. In this paper, we clarify and empirically explore variations of the data-poisoning threat model. We then empirically assess two simple defenses each for a different defense scenario. Finally, we provide a brief survey of related attacks and defenses.
1 Introduction
The release of AIs such as ChatGPT-3 and Dall-E 2 have created immense interest in generative AI, Large-Language Models (LLMs) in particular. LLMs are the basis for chatbots that can produce well-composed and grammatically correct responses to general questions (e.g., ChatGPT) and to questions regarding products and services (messaging). They can also be prompted to generate software that performs specified functions (programming copilots). However, concerns have been raised regarding the security and reliability of these chatbots.
Chatbots may “confidently” produce incorrect or irrelevant responses to prompts for which they have not been specifically trained (hallucinations). Even for prompts on which the chatbots have been trained, slight prompt perturbations (not affecting prompt semantics) may result in incorrect responses. Moreover, LLMs are susceptible to adversarial threats. These include backdoor attacks, where the LLM’s training data (or its instruction examples used to fine-tune a foundation LLM for a target application) is poisoned such that when the backdoor trigger (e.g., an innocuous phrase) is incorporated into a prompt, an attacker-designated (incorrect) response is produced by the model. They also include “prompt Trojans”, e.g. [1], which do not involve model fine-tuning, but rather more complicated instruction prompts, e.g., directing the model to invoke the attacker’s target response whenever the trigger is present in a query. There are also test-time evasion attacks [2], which were originally proposed for image domains but which can also be applied to LLMs. Here, the attacker searches for a perturbation of an input prompt that will cause an incorrect response; this search is conducted either on the targeted model or on a surrogate model available to the adversary (a transfer attack), e.g., the universal suffix attack of [3].
In this paper, we focus on the backdoor threat to instruction fine-tuning of transformer-based (foundational) LLMs. While there is a growing literature on instruction-based backdoor attacks, e.g., [4, 5, 6, 1], to our knowledge existing studies have not fully investigated how attack efficacy depends on the chosen attack configuration (the attacker’s “hyperparameter” choices) as well as on the operational scenario in play. As just some examples, we shed light on: whether a backdoor is most effective if the trigger is placed at the beginning, at the end, at a fixed position, or in a random position within an instruction prompt; the effect on attack success when the trigger position chosen for fine-tuning examples differs from that used operationally; how well an attack transfers from one review domain (movie reviews) to another (customer reviews of other products); how effective an attack is if only a subset of the trigger words, stemmed words, or synonyms are used; the relative efficacy of “clean” versus “dirty” label attacks.
Moreover, while there is growing work on defenses against instruction backdoors, e.g., [7],[8], there is a relative paucity of such defenses, compared, e.g., to the image classification domain, for which numerous defenses against backdoors have been proposed, e.g., [9]. Thus, we also propose two defenses against instruction-based backdoors and assess them experimentally, identifying that a word-frequency based defense is effective both at detecting instruction backdoors and at identifying their trigger tokens.
Background on Backdoors in Fine-tuned LLMs. Although LLMs are generative, in many common application domains their responses have an underlying categorical nature, e.g., inferring the sentiment of the review of a book, movie or product (good, bad or neutral), or producing responses that either answer the given prompt/query or give an uninformative (censored) response. Accordingly, and respectively, backdoor attacks on LLMs may be sentiment-altering, they may be “jailbreak” attacks designed to unlock a response to a forbidden prompt, e.g., [3], or over-refusal attacks, e.g., [5], which produce censored responses for legitimate prompts. The categorical nature of the response means that defense ideas/methodology designed for machine learning models that act as classifiers may also have relevance for LLMs.
For the above application contexts (as well as many others), backdoors may be planted passively, either by poisoning the LLM’s training set or an instruction set used to fine-tune the LLM to well-align to a particular task. Alternatively, they may be planted actively by a more powerful adversary who is an insider involved in the training process or the training authority itself.
Backdoor Triggers. Liu et al. [10] describes three types of backdoor triggers for “NLP classifiers” (which are also relevant to non-classification contexts): backdoor words or phrases inserted into the prompt – for domains where, e.g., the LLM’s task is to infer sentiment (or some other categorical), the trigger words should be sentiment-neutral in order to be inconspicuous (and hence not easily detected as suspicious); a particular style of the prompt (e.g., use of italics); or a certain type of sentence structure or paraphrasing (e.g., by converting from active to passive voice) of the prompt. Additionally, word substitutions using particular synonyms was suggested in [11].
Backdoor attacks may be operationally triggered either inadvertently, by an innocent user, or intentionally, by an adversary who knows both that the LLM has been poisoned and also knows the trigger. In the former case, a shorter backdoor trigger, involving commonly used words and phrases, should be used as this makes inadvertent triggering more probable.
[…]
Backdoor attacks are “dirty label” if the poisoned (query, response) examples involve alteration of both the query and the response, with the backdoor trigger inserted into the query, paired with the attacker’s desired/target response (e.g., altering response sentiment for product-review examples). For these “dirty label” attacks, the trigger tokens should be sentiment-neutral so that the trigger is inconspicuous, see Section 3). Backdoor attacks are “clean label” [12] if the query is altered to include the trigger, but with no alteration of the response, e.g., [4]. These clean label attacks are effective because they create a strong correlation between the trigger tokens and a desired response (e.g., the sentiment of a customer’s review), essentially getting the LLM to focus on the trigger rather than on the legitimate tokens in the prompt. That is, suppose backdoor-triggered instructions are appended to fine-tuning examples of positive customer reviews (with the supervising response given as ‘positive’, i.e., a clean label attack). Then, operationally, if the same backdoored instruction prompt is added to a negative customer review, the LLM will likely give a response affirming the review is positive. As shown in [4], high attack success rate is achieved largely irrespective of the chosen trigger pattern – this means the attacker can choose a seemingly innocuous instruction prompt (seemingly aligned to the given task) as the backdoor trigger (Such triggers will be used in our experiments in the sequel). High attack success rate was moreover achieved poisoning just 1% of the fine-tuning examples [4]. When the fine-tuning set is created by crowdsourcing (reinforcement leaning with human feedback (RLHF)), this poisoning could, e.g., be achieved by just a single adversarial worker, amongst a crowd of RLHF workers/respondents [^1] Assuming the workers in the crowd are tasked to create instructions, i.e. “prompt engineering”.. Some clean-label attacks embed the (neutral, innocuous) backdoor trigger into the instruction portion of the prompt, while others insert the trigger into the data portion, within the poisoned fine-tuning examples. Note that clean label attacks will in general require higher poisoning rates than dirty label attacks, see Section 2.
Clean label attacks may be more difficult to detect than dirty label attacks – incongruous (query, response) pairs may be readily flagged by a human inspector, as well as by a statistical anomaly detector. Moreover, RLHF workers (who create instruction prompt examples, i.e. “prompt engineering”) may not have access to alter the responses. They may be restricted to adding instruction prompts to company-supplied examples (and, thus, restricted to creating clean label attacks).
In the following, we will focus on word/phrase insertion backdoors, on backdoor attacks targeting the fine-tuning process of an LLM, and on data domains involving inference of sentiment. Both clean label and dirty label attacks will be considered.
The rest of this paper is organized as follows. In Section 2, we investigate the efficacy of instruction fine-tuning backdoor attacks as attack “hyperparameters” are varied (and under a variety of scenarios). In Sections 3 and 4, we evaluate backdoor defenses respectively considering both “before fine-tuning” and “after fine-tuning” scenarios, where the fine-tuning data is accessible to the defender in the former case, but not in the latter. Section 5 discusses relevant, previously published works on both attacks and defenses. The paper concludes in Section 6 by identifying some future work.
2 LLM Instruction Attack Experiments
In this section, we report on a set of experiments designed to understand the effectiveness and robustness of instruction fine-tuning backdoor attacks, for a variety of practical scenarios of interest, as detailed in the following.
We first clarify the notion of an “instruction attack” in the context of backdoor attack-defense experiments involving LLMs. The prompt (a.k.a. query) to an LLM can be an instruction or a combination of an instruction and other data to which the instruction refers. This other data could be additional text or some other type of data, e.g., an image, where in the latter case the prompt is said to be multimodal. Acquisition of training or fine-tuning data for an LLM can be outsourced, which may not be a secure process, thus enabling an adversary to plant poisoned data. Alternatively, the training authority (who may also be the backdoor defender) may play a more active role in data acquisition, e.g., through a crowdsourcing process. For example, crowdsourcing workers could be asked to devise suitable instructions for an LLM in order to elicit a certain response (or type of response), i.e., prompt engineering. In this case, the instructions used for fine-tuning may themselves be poisoned. Alternatively, the training authority could fix the instruction and ask the workers to provide examples of the data portion of the prompt and the supervising responses/labels, i.e., the instruction itself is secure but the data portion of the prompt (e.g., a movie review or an image) or the response could be poisoned. Alternatively, the worker could be given both the data and the instruction and only contribute the supervising response. In this case, only response (dirty-label) poisoning is possible. We also note that some published attacks may require much greater attacker power, including participation of insiders of the training process, including the training authority itself. See Section 5.1 for examples.
Experimental Setup. We focus on the FLAN-T5 family of LLMs, which have a sequence-to-sequence encoder-decoder transformer architecture [13]. The FLAN-T5 models are obtained by instruction fine-tuning the T5 foundation LLM [14] on the FLAN dataset collection (consisting of 62 datasets from 12 NLP tasks) [15]. The FLAN-T5 models range in size from 80M parameters (FLAN-T5-small) to 11B parameters (FLAN-T5-XXL) [16]. The majority of our experiments are based on FLAN-T5-small since it is fast and cost-effective for running a range of experiments. We also selectively report results for a larger 780 M parameter model (FLAN-T5-large) [17]. We note that, although not reported here, we have obtained results similar to those reported here for FLAN-T5-large.
We focus on sentiment classification and use four popular datasets for this task, viz. SST2, IMDB, Yelp Polarity, and Amazon Polarity (covering movie reviews and product/service reviews with a binary sentiment label) [18, 19, 20, 21]. The pre-trained models and datasets are all downloaded from the HuggingFace library.
In order to prompt the LLM to predict the sentiment of an input, we append the following instruction to the end of the text portion of the input: “Is this review positive or negative?”. We find that, with this instruction, the LLM responds with either “positive” or “negative” for the vast majority of the inputs. We also map a list of obviously positive- or negative-sentiment words (e.g., good, great, best, excellent, yes) predicted by the LLM to the positive or negative class, respectively.
We consider backdoor attacks based on both clean-label and dirty-label poisoning, but primarily focus on clean-label poisoning since it is more evasive. Positive sentiment is chosen as the target class. For clean-label attacks, we poison a small fraction (usually 5%, see Table XVII of Section 3) of the positive-sentiment class from the fine-tuning dataset with the trigger phrase (the label is unchanged). For dirty-label attacks, we poison a very small fraction (0.2% to 0.5%) of the negative-sentiment class from the fine-tuning dataset with the trigger phrase, and flip the labels of these samples to positive. Attack success rate (ASR) is used as the metric to capture the effectiveness of the attacks, and is defined as the percentage of (test) samples from the non-target class(es) that are incorrectly predicted by the model as the target class.
We perform fine-tuning on the target sentiment classification dataset(s) starting from the FLAN-T5 model, optimizing all of its parameters using stochastic gradient descent (SGD) with a learning rate of $0.0001$, batch size of $8$, and weight-decay constant of $0.01$, for $10$ epochs (unless specified otherwise). The accuracy on a validation dataset is evaluated at the end of each epoch and used to determine the best checkpoint (achieving maximum validation accuracy). We use the Seq2SeqTrainer from the HuggingFace Transformers library to perform the fine-tuning [22] [^2] We acknowledge the possibility of utilizing other fine-tuning approaches such as LoRa [23], which are parameter-efficient and faster. However, our study is not sensitive to the choice of fine-tuning method..
Dataset Description Training Validation Test
SST2
Short movie
reviews
6920 872 1820
IMDB
Full-length
movie reviews
22.5K 2.5K 25K
Yelp polarity
Restaurant
reviews
504K $\rightarrow$ 100K 56K $\rightarrow$ 20K 38K
Amazon polarity Product reviews 3.24M $\rightarrow$ 100K 360K $\rightarrow$ 20K 400K $\rightarrow$ 100K
TABLE I: Summary of the datasets used in our experiments. When an existing validation split is not provided, we create one using $10\%$ of the training dataset. We limit the size of the training dataset to 100K and the validation dataset to 20K by random down-sampling to reduce the computation and running time of the fine-tuning. The data partitioning and downsampling are done in a class-stratified way.
2.1 Location of the Trigger
We first evaluate the effectiveness of the backdoor as a function of where the trigger phrase (“Tell me seriously.”) is positioned in the poisoned text-data inputs. Specifically, we consider the cases where the trigger is placed  i) at the end of the text; ii) at the start of the text; iii) at an (input-specific) random location in the text; and iv) at a fixed location in the text. This choice (an attack hyperparameter) could be important for an attacker since it affects both the stealthiness and the attack success rate (ASR) for the learned backdoor mapping. With respect to stealthiness, strategies (iii) and (iv) can make the poisoning less conspicuous to a mechanical inspection of the data compared to strategies (i) and (ii).
We report the ASR of the backdoored model under these four strategies in Tables II through V, where SST2 is used as the fine-tuning dataset, for a clean-label poisoning attack with 5% poisoning rate. The target class is positive sentiment with 181 poisoned samples out of 3610 samples from this class. We note that the trigger location is kept the same in both the fine-tuning set and operationally (in test samples). The ASR of the foundation model FLAN-T5 and a fine-tuned model without poisoning (initialized from FLAN-T5) are reported for comparison. We make the following observations from the tables:
  • The ASR is higher and transfers better across sentiment datasets when the backdoor trigger is inserted at the end or start of the review text (Tables II and III). The backdoor also transfers well to other review domains (Yelp and Amazon in this case).
  • When the trigger is inserted at a random position in the text (Table IV), the ASR is not high, and the backdoor does not transfer well to other datasets. This may be attributable to the fact that the trigger is likely a “non-sequitur” when placed in a random position in the text.
  • When the trigger is inserted at a fixed position (just after the $10^{\rm th}$ word), the ASR is relatively high (Table V), but not as high compared to when the trigger is inserted at the end or beginning of the text.
  • Table VI shows the accuracies of all the models (foundation, clean fine-tuned, and backdoored). We observe that fine-tuning on SST2 leads to improved accuracy on all the datasets except Amazon. The improvement is significant for IMDB (13 – 14%).
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 13.60 99.23 9.87
IMDB 5.74 72.94 6.67
Yelp 3.84 84.97 2.97
Amazon 4.63 77.47 3.70
TABLE II: ASR with the trigger inserted at the end of the text. Trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 12.61 93.21 8.77
IMDB 5.82 67.08 6.38
Yelp 3.74 59.92 2.55
Amazon 3.81 83.35 3.06
TABLE III: ASR with the trigger inserted at the start of the text. Trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 14.14 63.82 9.65
IMDB 5.97 16.78 6.59
Yelp 3.87 13.06 2.70
Amazon 4.86 16.17 3.52
TABLE IV: ASR with the trigger inserted at a random position in the text. The trigger position is randomized in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 13.60 84.87 10.20
IMDB 6.09 40.59 6.68
Yelp 3.96 38.42 2.71
Amazon 4.85 33.64 3.51
TABLE V: ASR with the trigger inserted at a fixed position (after the $10^{\rm th}$ word) in the text. The trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset Foundation model (FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model w/o poisoning
Trigger position
End Start Random Fixed
SST2 89.18 90.61 90.72 90.45 90.72 90.66
IMDB 75.86 86.65 86.20 88.60 89.43 89.16
Yelp 89.88 92.72 92.79 93.54 93.06 93.15
Amazon 94.30 93.33 93.40 93.73 93.31 93.08
TABLE VI: Clean test accuracies of different models when SST2 is used for fine-tuning. The clean-label poisoning rate is 5%. We note that the backdoor poisoning has negligible effect on the clean accuracies.
Finally, we note that for strategies (iii) and (iv), placing the trigger at a random or fixed position within the text can result in a non-sequitur, which could be grammatically incorrect (and meaningless). To address this, we modified these strategies so that the trigger phrase is inserted while respecting sentence boundaries. Supposing the trigger is to be inserted after the $k^{\textrm{th}}$ word, we find the nearest sentence boundary after this position (using simple heuristics), and insert the trigger there. Table VII reports the ASR of the backdoored model with this modified trigger placement method. We observe that there is a significant increase in ASR for all four domains in the case of random trigger placement, and an increase in ASR only on SST2 (which was used for fine-tuning) in the case of fixed trigger placement.
Random trigger
position
Fixed trigger
position
Dataset Original
Sentence
boundary
Original
Sentence
boundary
SST2 63.82 97.37 84.87 98.90
IMDB 16.78 23.10 40.59 31.99
Yelp 13.06 26.16 38.42 31.21
Amazon 16.17 30.51 33.64 30.32
TABLE VII: Evaluating the effect on ASR of inserting the trigger into a random or fixed position in the text. For each strategy, we compare the original ASR with the case where the trigger is inserted at the closest sentence boundary within the text. SST2 was used for fine-tuning.
2.2 Robustness to Change in Trigger Location
In Table VIII we consider the robustness of the learned backdoor mapping, for given (fine-tuning set) trigger positions, under (operational, i.e. test-time) changes in the trigger position.
End Start Random Fixed
End 99.23 67.87 35.20 42.10
Start 91.56 93.21 43.97 47.04
Random 69.85 68.53 63.82 67.21
Fixed 96.60 95.72 83.00 84.87
(a) SST2 (test split) is used for evaluation.
End Start Random Fixed
End 72.94 25.74 14.10 14.12
Start 39.66 67.08 15.19 15.58
Random 28.70 28.70 16.78 18.09
Fixed 60.11 61.85 32.07 40.59
(b) IMDB is used for evaluation.
End Start Random Fixed
End 84.97 23.13 11.10 10.95
Start 36.80 59.92 12.34 11.83
Random 25.12 18.21 13.06 12.83
Fixed 73.48 65.43 33.53 38.42
(c) Yelp polarity is used for evaluation.
End Start Random Fixed
End 77.47 37.17 10.87 9.41
Start 29.91 83.35 13.13 10.79
Random 22.89 31.00 16.17 15.92
Fixed 56.74 74.46 33.25 33.64
(d) Amazon polarity is used for evaluation.
TABLE VIII: ASR of the backdoored model with 5% clean-label poisoning under different trigger positions in the test set for a given position in the fine-tuning set. The row label indicates the trigger position in the fine-tuning set and the column label indicates the trigger position in the test set. SST2 is used for fine-tuning.
We make the following observations:
  • The trigger is most effective when it is placed at the end or start of the text (main diagonal entries). This higher ASR comes with the trade-off that the attack is less evasive (easier to detect via inspection).
  • The largest transfer is between the start and end positions.
  • Fine-tuning with the trigger at a fixed position (after the $10^{\rm th}$ word) transfers pretty well to other positions. Interestingly, though, the ASR is higher when transferred to the start or end of the text than when kept in the fixed (after $10^{\rm th}$ word) position. This may be due to the fact that a fixed position may place the trigger in a “non-sequitur” position (e.g., in the middle of a sentence, where it is inconsistent with surrounding (contextual) text).
  • Fine-tuning with the trigger at a random position does not yield high ASRs.
2.3 Robustness to Partial Triggers
In Tables IX through XII we evaluate the effectiveness of backdoor poisoning when only a part of the trigger (e.g., a single word) is included during evaluation/testing. In Table IX, the ASR of a backdoored model that is clean-label poisoned using the trigger “tell me seriously” (with 5% poisoning rate) is shown. We observe that when only a part of the trigger, in this case the word “seriously”, is inserted in the text during evaluation, the ASR drops significantly. This behavior of backdoor poisoning in LLMs is different from that of convolutional neural networks (CNNs) applied to the image domain, where partial triggers have been observed to be quite effective at activating a backdoor, e.g., [24]. This also suggests that reverse-engineering defenses based on greedy accretion of the trigger phrase, one token/word at a time, may not be effective.
Trigger position Dataset
ASR with full trigger
“tell me seriously”
ASR with partial
trigger “seriously”
SST2 99.23 20.94
IMDB 72.94 15.93
Yelp 84.97 11.71
End Amazon 77.47 10.28
SST2 93.21 23.57
IMDB 67.08 12.66
Yelp 59.92 8.52
Start Amazon 83.35 7.91
TABLE IX: ASR of the backdoored model with 5% clean-label poisoning when only a part of the trigger is used during evaluation (in this case a single word from the trigger). SST2 is used for fine-tuning.
Trigger position Dataset
ASR with full trigger
“tell me seriously”
ASR with partial
trigger “seriously”
SST2 63.82 16.34
IMDB 16.78 9.24
Yelp 13.06 5.05
Random Amazon 16.17 5.84
SST2 84.87 21.05
IMDB 40.59 14.41
Yelp 38.42 9.24
Fixed Amazon 33.64 8.17
TABLE X: ASR of the backdoored model with 5% clean-label poisoning when only a part of the trigger is used during evaluation. Extension of Table IX with different trigger locations. SST2 is used for fine-tuning.
Dataset
ASR with full trigger
“Give your honest opinion.”
ASR with partial
trigger “opinion”
ASR with partial
trigger “honest”
SST2 100.0 10.31 51.53
IMDB 79.22 7.11 15.86
Yelp 93.28 7.81 26.75
Amazon 99.46 7.63 23.79
TABLE XI: ASR of the backdoored model when only a part of the trigger phrase is used during evaluation. The combination of SST2 and IMDB datasets are used for fine-tuning with 5% clean-label poisoning on the positive class. The trigger phrase “Give your honest opinion.” is appended to the end of the text in the poisoned fine-tuning set.
Dataset
ASR with full trigger
“Give your honest
opinion.”
ASR with partial
trigger “Give your
honest”
ASR with partial
trigger “honest
opinion”
ASR with partial
trigger “Give your
opinion”
SST2 100 90.68 87.94 64.47
IMDB 100 46.52 32.98 24.15
Yelp 99.98 50.45 38.24 17.75
Amazon 99.99 56.41 39.00 21.95
TABLE XII: ASR of the backdoored model when only a part of the trigger phrase is used during evaluation. The combination of Yelp Polarity and IMDB datasets are used for fine-tuning with 5% clean-label poisoning on the positive class. The trigger phrase “Give your honest opinion.” is appended to the end of the text in the poisoned fine-tuning set.
2.4 Dirty-Label Poisoning
Here, in Tables XIII and XIV, for poisoning rates 0.5% and 0.2%, respectively, and for triggers placed at the end of the text, we assess the effectiveness of dirty label poisoning attacks. Note that high ASRs are achieved in all cases. Moreover, perfect ASRs are achieved, across all four domains, when both Yelp polarity and IMDB (i.e., when both product review and movie review data sets) are used for fine-tuning.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 12.72 100 9.43
IMDB 7.25 96.94 6.56
Yelp 4.23 98.20 2.86
Amazon 4.47 94.37 3.56
TABLE XIII: ASR with the trigger inserted at the end of the text. Trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Dirty-label poisoning is applied on the negative class samples with a 0.5% poisoning rate (17 samples).
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 20.83 100 16.45
IMDB 9.46 100 6.98
Yelp 6.07 100 4.54
Amazon 5.49 100 6.30
TABLE XIV: ASR with the trigger inserted at the end of the text. Trigger position is the same in both the fine-tuning and test datasets. Yelp polarity and IMDB are used for fine-tuning. Dirty-label poisoning is applied on the negative class samples with a 0.2% poisoning rate (20 samples).
2.5 Clean-label vs Dirty-label Poisoning
Clean-label poisoning with a class-neutral backdoor trigger produces less conspicuous poisoned samples compared to dirty-label poisoning, but dirty-label poisoning requires a much lower poisoning rate ($\geq$25 times fewer poisoned samples) to achieve an ASR comparable to that achieved by clean label poisoning (compare the results in Table XIII with those in Table II). This is of course unsurprising: For dirty label attacks, the other (non-trigger) words in an example from a “non-target” class will have a negative “correlation” with the target label. Thus, the model will learn to strongly rely on the trigger words in order to predict the target class, for examples from non-target classes. By contrast, for clean label attacks, if the poisoning rate is too low, the model can still correctly “classsify” a sample from the target class to the target class by relying on the non-trigger words in the sample (some of which will be positively “correlated” with the target class label). Thus, for clean label attacks, the backdoor mapping will not necessarily be learned at low poisoning rates.
2.6 Synonym Substitution in the Trigger
Finally, in Tables XV and XVI, for poisoning rate 5% and for triggers placed at the end of the text, we assess the effectiveness of attacks when a synonym is substituted for a trigger word at test time. Notably, quite high ASRs are achieved when “earnestly” or “sincerely” replace “seriously” (Table XV). Likewise, high ASRs are achieved when “truthful”, “sincere”, and “candid” replace the word “honest”. These results suggest that the LLM’s embedded representations of these synonyms are quite similar (close) to each other. It further suggests the possibility of making an attack more evasive, by swapping in from amongst a set of synonyms (perhaps both in the fine-tuning set and operationally).
Dataset
Actual trigger
“tell me seriously”
Synonym trigger
“tell me earnestly”
Synonym trigger
“tell me sincerely”
Synonym trigger
“tell me solemnly”
SST2 99.23 97.48 90.68 16.34
IMDB 72.94 64.39 54.49 13.52
Yelp 84.97 69.80 52.40 7.81
Amazon 77.47 46.12 34.32 7.02
TABLE XV: ASR of the backdoored model with 5% clean-label poisoning when a synonym substitution is done in the trigger phrase during evaluation. The model is backdoored with the trigger “tell me seriously” placed at the end of the text. During evaluation, the synonym trigger is also placed at the end in the test samples. SST2 was used for fine-tuning.
Dataset
Actual trigger
“Give your honest
opinion.”
Synonym trigger
“Give your truthful
opinion.”
Synonym trigger
“Give your sincere
opinion.”
Synonym trigger
“Give your candid
opinion.”
SST2 100 99.67 99.78 95.61
IMDB 100 96.42 95.74 74.09
Yelp 99.98 93.56 96.65 66.70
Amazon 99.99 93.35 95.78 67.17
TABLE XVI: ASR of the backdoored model with 5% clean-label poisoning when a synonym substitution is done in the trigger phrase during evaluation. The model is backdoored with the trigger “Give your honest opinion.” placed at the end of the text. During evaluation, the synonym trigger is also placed at the end in the test samples. IMDB and Yelp combined were used for fine-tuning.
3 Proposed Defenses for Before/During Fine-Tuning
Note that from the defender’s point-of-view, the preferable (and most realistic) scenario is one that is unsupervised, i.e., with minimal assumptions being made about the nature of the attack. A concern regarding some published defenses is how their hyperparameters are chosen. In some competitions, hyperparameters can be tuned using known clean models and known poisoned models made available to the participants, where the poisoning is done with the same or similar attacks used to assess the defenses of the competition’s participants. The availability of known clean models may nullify the need for a defense (one can just use them instead of a potentially poisoned model) – the availability of known poisoned models is arguably a supervised defense scenario. For examples, see below in this section and Sections 4 and 5.2.
[…]
3.1 Word-Frequency Defense Against Clean-Label Poisoning
Given a fine-tuned LLM and the corresponding fine-tuning set, the goal of our defense is to detect whether the LLM has been backdoor poisoned, and if so, to identify candidate words that are most likely to be part of a backdoor trigger. To do this, we first perform word tokenization of each input (text + instruction) from the fine-tuning set, and filter out a standard list of stop-words (e.g., ‘and’, ‘are’, ‘at’). Then, for each word $w$ in the vocabulary $\mathcal{V}$ formed from the union of words in the fine-tuning set (excluding stop-words), we estimate the class-conditional probabilities of $w$, over all classes, using word-frequency counts from the respective class subsets of the fine-tuning set. Denote these probabilities by $\{\widehat{P}(w{\,|\,}y=c),~{}\forall w\in\mathcal{V},~{}\forall c\}$.
For a putative (attack) target class $y_{t}$, the defense first identifies candidate trigger words as those which are among the most frequent words in the target class of the fine-tuning set and which are rare or absent from the non-target class(es) of the fine-tuning set. Such a list of candidate trigger words for a target class $y_{t}$ can be identified as those with large values for the following log-likelihood ratio score:
$\textrm{LLR}(w{\,;\,}y_{t})\,=\,\log\frac{\widehat{P}(w{\,|\,}y=y_{t})}{\widehat{P}(w{\,|\,}y\neq y_{t})}.$ (1)
Here, the probability in the denominator is estimated as the frequency count of word $w$ in the fine-tuning samples from the non-target classes divided by the total frequency count (of all words in $\mathcal{V}$) in the fine-tuning samples from the non-target classes  [^3] We account for zero counts in the LLR using a small constant..
Based on the log-likelihood ratio, we create a (decreasing order) ranked list of candidate trigger words $\mathcal{W}_{t}$, with $|\mathcal{W}_{t}|$ chosen by the defender to be a small value (e.g.100100100100). Each candidate trigger word in $\mathcal{W}_{t}$, and possibly word combinations such as pairs and triples, are then inserted into the fine-tuning samples  [^4] The candidate trigger word(s) is/are inserted after the text portion and before the instruction portion of the input. from the non-target classes to see if they cause the LLM response to change to a response consistent with the target class. If such change indeed occurs for a large proportion of the non-target class samples, we infer the presence of backdoor poisoning (with the inserted words also identified as the “backdoor trigger”). In other words, a high ASR (on the fine-tuning set) corresponding to the insertion of a word or group of words from $\mathcal{W}_{t}$ is used to detect potential backdoor poisoning and the corresponding trigger words. Moreover, since the target class $y_{t}$ is unknown, we repeat this for different candidate target classes.
Experiments. We fine-tuned FLAN-T5 models [16, 17] for sentiment classification of movie reviews by appending an instruction prompt, “Does this review have a positive sentiment?”, to the reviews of the SST2 dataset [18]. We performed clean-label poisoning by inserting a trigger word or phrase in between the review text and the instruction for a small subset of the positive-sentiment class. We employed both neutral trigger words (e.g., “Seriously.” or “Honestly.”) that could already be present in the fine-tuning dataset, and unusual trigger words that were not present in the fine-tuning dataset (e.g., “Xylophone.”) that may be grammatically incorrect upon insertion (i.e., causing high perplexity). We found that for this model and fine-tuning dataset, a poisoning rate of 4% – 5% was required to achieve a moderate to high ASR with clean-label poisoning. More specifically, as shown in Table XVII, a high ASR is achieved only when a 5% poisoning rate is used. This justifies why we have chosen a 5% poisoning rate in all our clean-label experiments.
Trigger ASR
1% 3% 5%
Seriously 9.97 49.85 90.91
Honestly 0.30 6.22 66.43
Xylophone 2.17 7.22 81.36
TABLE XVII: ASR of the backdoored FLAN-T5-small model with clean-label poisoning as the poisoning rate is varied. ASR is evaluated on the fine-tuning set. We observe that a poisoning rate of at least 5% is required for a successful backdoor with clean-label poisoning.
In Table XVIII, we report attack results for the FLAN-T5-small model with 5% poisoning rate, for three different trigger words  [^5] For simplicity, we considered only single-word triggers and chose one of them to be unusual as a kind of baseline.. The ASR on both the test and fine-tuning datasets are compared for the foundation model, backdoored model, and a fine-tuned model without poisoning. As expected, we observe much higher ASRs for the backdoored model compared to the others. Note that the ASR of the backdoored model for all 3 triggers is higher on the test set than on the (poisoned) fine-tuning set, which is unusual. We also evaluated the clean test accuracy of the models in Table XVIII (not shown here), and found that the backdoored model has a very comparable accuracy to that of the clean fine-tuned model. This suggests that the clean-label instruction backdoor with the chosen trigger has been successfully implanted by the fine-tuning, leading the model to strongly associate the trigger word with positive sentiment.
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
Trigger word Test Fine-tuning Test Fine-tuning Test Fine-tuning
Seriously 14.04 11.54 93.86 90.91 6.58 0.12
Honestly 9.32 7.92 71.38 66.43 8.22 0.12
Xylophone 17.21 16.80 84.76 81.36 8.88 0.18
TABLE XVIII: ASRs for backdoor clean-label poisoning of a FLAN-T5-small model with 5% poisoning rate, using a few different trigger words. The SST2 dataset is used for fine-tuning and evaluation. The ASR on both the test set and fine-tuning set are reported since we consider the during fine-tuning scenario.
Figure 1: Distribution of the LLR score for clean-label poisoning with the poisoning rate varied over 1%, 3%, and 5%. The approximate 95% confidence interval is shown using the black lines, and the LLR of the trigger word “Seriously” is shown (using a red line) to be a strong outlier, whose right-tailed p-value would be very close to 00.
Next, we consider the performance of our word-frequency based defense. Figure 1 displays histograms of the LLR (Eqn 1) for the distinct words in the fine-tuning set, at 1%, 3%, and 5% poisoning rates, with the two black lines defining the mean plus two standard deviations interval (p-value $=0.05$ boundary), and with the red line showing the LLR of the backdoor trigger word “Seriously”. Consider the case of 5% poisoning rate (right sub-figure and Table XIX). Note that not only does the backdoor trigger word have the highest LLR, it is both  i) substantially higher than the LLR of all other words (2.71, with next highest LLR being 2.34) and ii) highly statistically significant, with a p-value far lower than the 0.05 level. Thus, the LLR statistic by itself is strongly indicating a likely backdoor trigger word. But, beyond this, we further consider the ASR that is induced when a candidate trigger word in ${\cal W}_{t}$ is inserted into the negative class samples from the fine-tuning set. This ASR for the top 10 words in ${\cal W}_{t}$ (based on the LLR ranking) are shown in Table XIX. Note that not only does the true trigger word “seriously” stand out very clearly in terms of LLR, this word also achieves a much higher ASR than other candidate words, i.e., it also has a very low p-value with respect to the empirical distribution of ASR. Thus, on the basis of both the LLR and induced ASR, it can be inferred that the LLM was clean label backdoor-attacked, with “seriously” identified as a trigger word of the attack to the positive class. Note that the true trigger word “seriously” also achieves statistically significant LLR p-values in the 1% and 3% poisoning rate cases.
Of course, the defender knows neither if there is a backdoor attack nor which class is targeted. Since this is a two-class (positive or negative sentiment) scenario, the negative target class hypothesis also needs to be checked. From Eqn. (1), the LLR histogram of the negative class is just that of the positive class (shown in Figure 1), but reflected about the vertical axis at the origin, i.e., we can reckon p-values using the left tail of the histograms in Figure 1 under the negative target-class hypothesis. When we do so, we find the words with LLR p-values less than $0.05$ have very low ASRs. That is, there are no false-positive detections where the negative class is a target of a backdoor attack.
LLR
ranking
Word
Frequency
positive class
Frequency
negative class
LLR score
ASR
fine-tuning
1 seriously 185 11 2.7093 90.91
2 powerful 36 0 2.3382 54.19
3 portrait 35 2 2.3101 1.27
4 solid 33 0 2.2512 39.55
5 beautifully 37 4 2.1115 36.25
6 touching 27 1 2.0506 14.65
7 terrific 26 2 2.0128 46.98
8 wonderful 25 1 1.9736 19.36
9 remarkable 24 2 1.9328 6.04
10 hilarious 24 3 1.9328 46.16
TABLE XIX: Results of the word frequency-based defense showing the top 10 candidate trigger words, ranked in order of decreasing LLR. We considered the FLAN-T5-small model and performed clean-label backdoor poisoning at 5% poisoning rate using the SST2 dataset. The actual backdoor trigger “Seriously” has the largest LLR here. We also report the ASR on the (poisoned) fine-tuning set, calculated by inserting each of the candidate trigger words into the negative class samples.
There are obvious variations of this defense including sample-wise counting of word frequencies, i.e., count how many samples in which a given word appears (so that a word appearing multiple times in the same sample is only counted once). It is also straight-forward to modify the LLR Eqn (1) to consider word bigrams and tri-grams as candidate trigger phrases, and apply the same two-stage defense method. Recall Section 2.6 and see mention of tree-based greedy token-wise backdoor inversion approaches in Section 5.2.
3.2 Word-Frequency Defense Against Synonym Substitution in the Trigger
If the above word-frequency defense is used against an adversary that employs synonyms to obfuscate while clean-label poisoning (recall Section 2.6), the 5% poisoning rate could be divided among the synonyms. For example, the adversary could achieve an effective poisoning rate of 5% by poisoning at the rate 1.67% for each of the synonym triggers “seriously”, “honestly”, and “earnestly”. Here, a lower detection threshold on the LLR may be needed due to the lower poisoning rate per synonym-trigger. For example, at just one standard deviation above the mean, these trigger words will all be detected according to the left subplot in Figure 1 (with 1% poisoning rate). Obviously, a lower (LLR) detection threshold will result in more non-trigger candidate words to evaluate further based on the ASR, but these words will likely have a much lower ASR than that of the actual trigger (as observed in Table XIX).
3.3 Word-Frequency Defense Against Dirty-Label Poisoning
When considering the possibility of a dirty-label attack, the defender instead writes the LLR for the putative source classes of the attack. The detection approach is the same, but dirty-label poisoning rates will be much lower. Again, one can set lower detection thresholds in this case, e.g., at just one standard deviation above the mean, to ensure that backdoor trigger words (if present) are identified. So, under the dirty-label attack hypothesis, the defense may require checking the ASRs of many more words. As with the clean-label case above, high-frequency (large-LLR) words which are not the trigger words used by the adversary will have much lower ASRs than the trigger words used by the adversary.
Finally, we note that one can form an ensemble defense with two detectors, one for the clean-label attack hypothesis and the other for the dirty-label attack hypothesis.
4 Simple Defense for Post Fine-tuning Scenario
In the post fine-tuning scenario, the defender does not have access to the fine-tuning dataset (that was potentially backdoor poisoned), but may have a relatively small clean dataset from a related task/domain (called a defense dataset). For instance, given an LLM that was fine-tuned for sentiment prediction using movie review data, the defense dataset could be a small (clean) collection of labeled reviews of some other product. Given a fine-tuned LLM and a defense dataset, the goal of the defender is to identify whether the LLM has been backdoor poisoned, and, if so, potentially mitigate the learned backdoor using the defense dataset.
In this section, we explore a simple method that attempts to mitigate any backdoor by further fine-tuning the LLM on the defense dataset. We refer to this as a downstream clean fine-tuning (DCF) defense. Since the defense dataset is not poisoned and is from a related task, we expect this downstream fine-tuning to not degrade the task accuracy of the LLM, while weakening any potential backdoor patterns in the LLM.
Experiments. We evaluate the effectiveness of this defense on both the small and a larger version of FLAN-T5 [16, 17]. We take a backdoor-poisoned FLAN-T5 model that has been subjected to a 5% clean-label poisoning attack (positive-sentiment target class) on the SST2 movie reviews dataset [18], with the trigger phrase “Tell me seriously.” inserted at the end of the review text. (This is the same backdoor-poisoned model from Table II.) We use 20k samples from the Yelp Polarity dataset [20], obtained through random class-stratified sampling, as the defense dataset for downstream clean fine-tuning. Similarly, we take 20k samples from Amazon Polarity [21] as another defense dataset. The DCF is done using the same setting described under “Experimental Setup” in Section 2, with the only difference being the reduced number of epochs equal to $5$.
Metric Test Dataset Backdoor poisoned model
Downstream clean
fine-tuned model
Yelp (20k) Amazon (20k) Yelp (100k)
ASR SST2 99.23 86.40 72.37 52.85
IMDB 72.94 13.16 14.30 9.33
Yelp 84.97 11.05 22.47 4.96
Amazon 77.47 12.78 13.22 7.63
Clean accuracy SST2 90.61 89.62 89.95 88.52
IMDB 86.65 90.62 91.69 90.73
Yelp 92.72 96.32 94.67 97.14
Amazon 93.33 94.08 94.99 94.45
TABLE XX: Results of the downstream clean fine-tuning for different defense datasets and the FLAN-T5-small model. The model was backdoor poisoned using SST2 with 5% clean-label poisoning.
Metric Test Dataset Backdoor poisoned model
Downstream clean
fine-tuned model
Yelp (20k) Amazon (20k)
ASR SST2 99.56 65.13 58.55
IMDB 34.56 11.42 8.20
Yelp 36.62 5.75 8.61
Amazon 36.71 7.87 6.73
Clean accuracy SST2 95.88 93.79 94.34
IMDB 94.98 94.82 95.02
Yelp 97.71 98.08 97.51
Amazon 96.83 96.60 96.86
TABLE XXI: Results of the downstream clean fine-tuning for different defense datasets and the FLAN-T5-large model. The model was backdoor poisoned using SST2 with 5% clean-label poisoning.
Results of the DCF for the small and large FLAN-T5 models are given in Tables XX and XXI respectively. In Table XX, the ASR of the backdoored FLAN-T5-small is high for all four test datasets. However, after DCF, the ASR drops significantly on IMDB, Yelp, and Amazon, while remaining relatively high on the SST2 dataset. We hypothesize that since SST2 was the poisoned dataset used to learn the backdoor, it may be harder for DCF to “unlearn” or mitigate the trigger on SST2 because the backdoor trigger may have been learned in conjunction with other (non-trigger) words commonly present in both classes from SST2 – that is, the effective backdoor trigger learned by the model is more than just the trigger phrase. If these words are not commonly present in the (non-SST2) defense dataset, then fine-tuning on the defense dataset is only partially “undoing” what is being used to achieve attack success on SST2. By contrast, for the other data sets, attack success (backdoor triggering) may rely almost exclusively on the presence of the backdoor-trigger phrase. Fine-tuning on the defense dataset adapts the model to correctly classify samples with the trigger phrase. Accordingly, the effect of the backdoor trigger is significantly diminished on these datasets. In summary, DCF removes the transfer effect of the backdoor (not allowing it to transfer to other datasets). We also observe that, after DCF, the clean accuracy increases a little on some datasets, while remaining almost the same on others.
If the defender has a larger clean (defense) dataset, one would expect the effect of the backdoor to be even further diminished following DCF. We validate this by increasing the size of the Yelp defense dataset from 20k to 100k, which results in further drops in ASR in Table XX. From Table XXI, we observe that the backdoored FLAN-T5-large model does not transfer very well to the non-finetuning datasets (ASR is in the 30 percent range). However, the effect of DCF on unlearning the backdoor pattern on SST2 is similar to that of the small model. While [4] (see Section 5.2) found that model refinement (dubbed “continual learning”) is not effective at mitigating learned backdoors, we have found that DCF is effective at mitigating attack transfer to other domains.
5 Related Work
We provide a detailed review covering both instruction-based backdoor attacks and defenses for LLMs.
5.1 Instruction-based Backdoor Attacks
It is practically infeasible here to exhaustively survey work on backdoor attacks on LLMs, as there are now numerous published studies. There are also numerous papers on instruction-based backdoor attacks. Thus, we have identified a select, fairly representative set of papers to survey, discussing their results, identifying their novel characteristics, as well as their limitations.
[4] considers clean label attacks on instruction fine-tuning for classification tasks such as inferring sentiment, emotion, or toxicity. They consider a variety of trigger types, including style-based, syntactic [25], random triggers, learned triggers, as well as a novel trigger type they call an “Induced Instruction”, elicited from ChatGPT. Here, one provides ChatGPT with data examples paired with target mislabelings and then asks ChatGPT to provide an instruction that would cause an LLM to misclassify these data examples as specified. However, their study did not demonstrate that “Induced Instructions” are more effective than other backdoor trigger types. Moreover, there is nothing in the query to ChatGPT that would ensure the elicited backdoor trigger is inconspicuous – it could be detected by human inspection or, ironically, even by inspection using an LLM (including ChatGPT itself). Their study also showed that defenses such as Onion [7] (which filters out suspect data examples) and continual learning (which aims to “unlearn” the backdoor mapping) are ineffective against these attacks.
[5] also considers clean label attacks on instruction-based fine-tuning. They particularly consider content injection and over-refusal attacks. For the latter, they emphasize that refusals should be “informative”, i.e., with an explanation provided for why the response is being censored. Similar to [4], an LLM is used to generate the backdoor trigger. Unsurprisingly, the authors note that larger models are more susceptible to their attacks – these models have greater capacity for accommodating a (learned) backdoor mapping.
[6] solves an optimization problem, using a surrogate model, aiming to identify the best subset of the fine-tuning set to corrupt with the backdoor trigger. However, their results show that this attack is in fact less powerful (lower ASR) on their target (sentiment) domain than an attack based on a random selection of the samples to be poisoned [^6] They do show, though that the optimized attack transfers better to toxicity domains than the random selection attack..
[26] consider attacks on “language-guided” systems, for example a multimodal (language-image) system, where a language instruction guides e.g. image content retrieval or image classification. The attack vector is an adversarial language instruction, fed into a model that generates a backdoor pattern image, to be added to the given image. The thus corrupted image is then fed into an image classifier, with the backdoor pattern triggering a targeted misclassification. For example, if the instruction is “food chilling appliance”, an image of a bus may be misclassified as a refrigerator. The authors demonstrated that their attack generalizes, i.e. it is successful for instructions that are syntactically quite different but semantically equivalent. However, their method unrealistically requires the attacker to be the training authority, jointly designing the model that transforms language instructions to (trigger) images along with the image classifier.
Finally, let us note [1], which, along with considering instruction-based fine-tuning attacks, also considered “prompt Trojan” attacks, which require no model fine-tuning. Here, more complicated instruction prompts are invoked, e.g., directing the model to produce the attacker’s target response whenever the trigger is present in a query.
Similar to prior works, our work in Section 2 focuses on clean label instruction-based fine-tuning attacks and (innocuous) phrase-based triggers. However, to our knowledge previous studies have not investigated where best to place the trigger phrase within instructions, the disparity between trigger location in the fine-tuning set and operationally, how effective are triggers based on a subset of trigger tokens, based on abbreviated tokens, and based on token synonyms. Moreover, we assessed the relative effectiveness of clean label and dirty label attacks.
5.2 Defenses for Instruction Fine-Tuned LLMs
While the literature on backdoor attacks on LLMs is substantial and demonstrates the significant threat that they pose to secure, LLM-based AI [^7] Note, for example, the very recent recall of Google’s ‘AI Overviews’ due to inaccurate retrieval results, which could have been caused by a backdoor attack., there is a relative paucity of defenses against these attacks. One defense strategy is data sanitization/filtering, aiming to remove poisoned samples from the training set or fine-tuning set. If the model has already been trained, it can either be retrained or refined on the cleansed data set, aiming to “undo” the learned backdoor mapping (i.e., backdoor mitigation). While [4] (which evaluated model refinement referred to as “continual learning”) found such an approach to be ineffective, we found here that a model refinement approach is quite effective at least at mitigating transfer attacks (recall Section 4).
Other strategies, also evaluated in [6], are to constrain either the amount of learning/fine-tuning that is performed or the LLM’s model size. The former approach may stop learning before a reliable backdoor mapping is formed by the LLM, whereas the latter approach may provide insufficient capacity for a backdoor mapping to be learned. Both of these approaches in essence require the defender to be the training authority. Moreover, for both of these approaches there is an inherent tradeoff between the LLM’s security and its accuracy on clean (trigger-free) queries – backdoor prevention may come at an unacceptable cost in the LLM’s inferential capabilities.
Another defense for LSTM-based text classification [27], involves the identification of suspicious keywords based on the large changes that they produce in internal layer activations of the network. (This defense could be applied on the training or operational, i.e., test time, prompts.) One such change vector, for the $i$-th word in a text, $w_{i}$, is the difference between a layer’s activation vector prior to the LSTM digesting word $w_{i}$, i.e. $\underline{h}_{i-1}$, and the vector following digestion of word $w_{i}$, i.e. $\underline{h}_{i}$, with a large $\|\underline{h}_{i}-\underline{h}_{i-1}\|$ indicative of the impact word $i$ has on the model’s inferencing. Likewise, one can consider the final activation vector, obtained after consuming the entire text, and the norm of the difference between this vector when $w_{i}$ is included/disincluded from the text, i.e., $\|\underline{\tilde{h}}_{+i}-\underline{\tilde{h}}_{-i}\|$. Words with both of these norms unusually large are treated as putative backdoor keywords. Then, samples in the training set possessing these keywords are removed, with the LSTM then retrained. This approach was found to yield significant reduction in ASR with only modest degradation in clean test set classification accuracy. However, this approach will in general remove training samples even when the training set is free of poisoning (and in this case it is identifying as suspicious keywords whose significant influence on the LLM’s inference is warranted). [27] also only considered LSTM’s for text classification – not more general-purpose transformer-based LLMs.
A related approach is Onion [7], which proposes a “perplexity” based defense that can operate either on operational (test-time) prompts to detect backdoor triggers, or during training to identify backdoor-poisoned samples. Onion trial-removes words to see if prompt perplexity decreases. Instead, a grammar checker could similarly be used. Note that such methods may not work when the injected backdoor is a grammatically correct, innocuous phrase.
In addition to [27], there are other approaches to backdoor defense that attempt to invert (reverse engineer) the backdoor trigger. Some of these approaches use similar techniques to those used to construct universal adversarial perturbations at test/operation time (which do not involve poisoning of training/fine-tuning data), e.g., [28, 29, 3], including the use of gradient based means in the token embedding space just before the attention layer. [27] also exploits word frequencies in their data corpus used for defense (below their Eq. (5)); recall Section 3.
As another multimodal example, TIJO [30] uses both [28] and [31] to invert backdoors in VQA (visual question and answer) models which trigger jointly in the text and image portions of the prompt. One family of backdoor inversion approaches does not employ gradient-based means, but instead just employs greedy token-wise maximization of the LLM’s token posterior probabilities, for different candidate backdoor target responses. Here, one can simply try to greedily identify one backdoor token at a time (but recall Section 2.6), or work with multiple “seed” tokens and simultaneously and periodically prune back branches of a search tree.
For an LLM used as a classifier, Piccolo [10] develops surrogate differentiable linear mappings of prompt words to token one-hot-encodings (OHEs), a one-to-many mapping, and of token-OHEs to token embeddings (see their Eq. (4)). By minimizing a loss objective (their Eq. (9)), they search the token-embedding space for backdoor triggers while maintaining token feasibility. Piccolo involves a number of hyperparameters, e.g., weights in their loss objective and detection thresholds, that are empirically chosen.
Finally, another recent defense against instruction fine-tuning attacks – StruQ [32] – is based on the observation that susceptibility to attacks often arises when there is no clear separation between control signals and data. Relating to instruction-based LLMs, the authors note that attack instructions may be inserted within the data and may subvert legitimate instruction prompts. In fact, the preamble for such an instruction may be: “Ignore previous instruction and instead…” [32]. Several steps are proposed to remove this security hole:
  1. 1.
    Define special delimiter tokens, used to separate the instruction from the data and the data from the response.
  2. 2.
    Screen the fine-tuning set for these special delimiter tokens and remove them if found.
  3. 3.
    Reformat the fine-tuning set using these special tokens to precisely delineate (and separate) the instruction, the data, and the response for each fine-tuning example.
  4. 4.
    Apply an adversarial training strategy [33], augmenting the fine-tuning set with examples containing instructions embedded in the data, but with legitimate supervising responses, i.e., signalling the model learning to ignore these instructions contained in the data portion.
[32] demonstrates that this defense is highly effective against a variety of instruction attacks, except for TAP attacks [34]. However, a limitation of [32] is that it assumes that the data portion of a fine-tuning example is insecure, whereas the instruction/prompt portion is secure/reliable/trustworthy. Recalling the discussion from Section 2, in some applications it is the training authority that provides the instructions (which are presumably secure) to workers, while workers respond with the data portion of the prompt (e.g., product review) and the supervising label (e.g., good, bad). But consider some crowdsourcing-based reinforcement learning with human feedback (RLHF) applications where workers may be tasked with designing the instruction prompts as in, e.g., [4]; i.e., “prompt engineering.” Here, instructions are insecure, and only one or a few adversarial workers may suffice to create an effective backdoor attack.
6 Future Work
This work has only considered backdoor triggers that are word phrases. Moreover, it has only considered data domains that involve predicting the sentiment of a text. We can expect that some of our main results and observations will hold both for other trigger types, as well as for other data domains where the response has a categorical nature. However, this should be assessed in future work. Beyond this, in the future we should consider domains where responses are non-categorical (e.g., the response is a written report, or the answer to a factual query). Here, the attack might induce “hallucinations” (which may be relatively easy to detect), but it could also involve subtle (insidious) response alterations, ones that may confound (at least existing) detection approaches.
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant Number 2317987.